​A basket of goods for a given consumer includes two goods, X and Z

A basket of goods for a given consumer includes two goods, X and Z. Consumer income is equal to $1,000 and the prices of these two goods are as follows: 

Px = $20 

Pz = $20 

This consumer is consuming 10 units of good X. 

Suppose that over the course of a year, the price of good X changes by - 10% and the price of good Z changes by 10%. 

How much income would be required for the consumer to afford the same quantity of goods X and Z with the new prices? $_______  

What is the rate of inflation? _______  % (Enter your response as a percentage rounded to two decimal places.) 

Given this change in prices, is it possible for our consumer to buy the original bundle of goods? _______ 
Answer #1

Income M = 1000

Spending on X = 10*20 = 200

Spending on Z = 1000 - 200 = 800

Quantity of Z = 800/20 = 40

Hence, current bundle is (10X, 40Z)

Price of X changes to 20 - 10%*20 = 18

Price of Z changes to 20 + 10%*20 = 22

Total spending on new bundle = 18*10 + 22*40 = 1060

Hence, income required for consumer to buy same bundle at new price = $1060

Rate of inflation = (1060 - 1000)*100/1000 = 6%

No. This is because to buy original bundle new income should be higher $1000.